solve the inequality: 1-2x greater than/equal to 5
Answer by arunpaul(104)

1-2x >= 5
or 1-2x-1 >= 5-1
or -2x >= 4
or 2x <= -4
or x <= -4/2 = -2
or x <= (-2) (Ans)